What is the maximum number of water molecules pr
Answer #1

17. from the data given we can do the following:

7.5gN_2O_3*\frac{1mol}{108g}*\frac{4molNO_2}{2molN_2O_3}*\frac{46gNO_2}{1molNO_2}= 6.4gNO_2

18. We can do the following taking into account that a stable molecule has an oxidation state = 0

Na = +1

ClO2 = -1

we can gladly help you with the others but you should post them again